Find the value of x so that the area under the normal curve between μ and x is approximately 0.4800 and the value of x is greater than μ.

1
Expert's answer
2021-12-06T11:25:37-0500
"P(Z<\dfrac{x-\mu}{\sigma})-0.5=0.4800"

"P(Z<\dfrac{x-\mu}{\sigma})=0.9800"

"\dfrac{x-\mu}{\sigma}\approx2.053749"

"x\approx\mu+2.053749\sigma"

Given "\mu=185, \sigma=18."


"x\approx185+2.053749(18)"

"x\approx221.967"